OpenStudy (anonymous):

A spherical raindrop evaporates at a rate proportional to ists surface area.Write a differential equation for the volume of raindrop as a function of time

OpenStudy (perl):

dV/dt = -kS , where S is surface area at time t

OpenStudy (perl):

S = 4Pi * r^2 V = 4/3 Pi * r^3

OpenStudy (perl):

so you can solve S in terms of V

OpenStudy (perl):

in V , solve for r V = 4/3 Pi * r^3 ((3/4)* V /Pi)^(1/3) = r

OpenStudy (perl):

now plug this into S , and you will have a differential equation in terms of V or t only

OpenStudy (anonymous):

V(t)=4*pi*((3/4)/V/pi)^(2/3) is that true

OpenStudy (perl):

The type of differential equation you are solving here has one dependent variable (time) and one independent variable (volume), and derivatives of the dependent variable with respect to the independent variable

OpenStudy (perl):

wait

OpenStudy (perl):

you need to plug that into S, not V

OpenStudy (perl):

V = 4/3 Pi * r^3 r = (3V/(4*Pi) ) ^(1/3) plug into S S = 4Pi * [(3V/(4Pi))^(1/3)]^2

OpenStudy (kainui):

Here check this out why not: \[\LARGE \frac{\partial V}{\partial t}=-k \frac{\partial V}{\partial r}\] Chain rule\[\LARGE \frac{\partial V}{\partial r}\frac{\partial r}{\partial t}=-k \frac{\partial V}{\partial r}\] divide out dV/dr\[\LARGE \frac{dr}{dt}=-k\] \[\LARGE r=-kt+C\] Now you just plug into the volume formula \[\LARGE V(t)=\frac{4}{3} \pi (-kt+C)^3\]

OpenStudy (perl):

now you have dV/dt = -k*S dV/dt = -k * 4Pi * [(3V/(4Pi))^(1/3)]^2

OpenStudy (alekos):

I concur

OpenStudy (alekos):

well done

OpenStudy (perl):

dV/dt = -k*S dV/dt = -k * 4Pi * [(3V/(4Pi))^(1/3)]^2 dV/dt = -k * 4Pi * [ (3V)/(4pi)] ^2/3 dV/dt = -k * 4Pi * [3/(4pi)] ^(2/3) * V^(2/3) dV/dt = - c * V^(2/3)

OpenStudy (kainui):

I think you're wasting time, \[\LARGE S=\frac{dV}{dr}\] as your substitution makes it easier to just apply the chain rule and solve.

ganeshie8 (ganeshie8):

S = V' makes sense right ?

ganeshie8 (ganeshie8):

with.respect.to r

OpenStudy (kainui):

Yeah @ganeshi8 and in this case our function looks explicitly like this: \[\LARGE V(r(t))\]

OpenStudy (perl):

looks like you solved the differential equation :)

OpenStudy (perl):

then you can take dV/dt , that is what the question asked, if i read correctly

OpenStudy (perl):

our answers should come out the same, after some substitution perhaps

OpenStudy (kainui):

@perl yeah I think the question is poorly worded, it's already a function of t, because r is a function of t. I think one of the trickiest things is that people aren't able to correctly write out what their function is, like perhaps they write z(t) or z(x,y) but really they have something like z(x(t),y(t))

OpenStudy (perl):

did you use partial diferentiation notation intentionaly would it be wrong to write dV/dt = -k* dV/dt , with just normal script d

OpenStudy (perl):

since you are not treating the other independent variable as a constant

OpenStudy (kainui):

To show that it works for this\[\LARGE \frac{d V}{d t}=-k \frac{d V}{d r}\] \[\LARGE V(t)=\frac{4}{3} \pi (-kt+c)^3 \\ \LARGE V'(t)=4\pi (-kt+c)^2 (-k)\] \[\LARGE V(r)=\frac{4}{3} \pi r^3 \\ \LARGE V'(r)=4 \pi r^2 \] \[\LARGE 4\pi (r)^2 (-k)=-k*(4\pi r^2)\] Yeah I guess I sorta just wrote them with partials for no reason, but it doesn't really matter since partial derivative of a single variable function is really the same thing. I don't know, I just felt like it and didn't think too much about that lol.

OpenStudy (perl):

ok just checking.

OpenStudy (perl):

i have seen people use partials with one variable , but this is like a composite function V(r(t))

OpenStudy (kainui):

Also there is not an "other independent variable" here, only t is independent, radius, surface area, and volume all depend on this variable only.

OpenStudy (perl):

ok

OpenStudy (perl):

i meant, V is a function of r, and r is a function of time. so you could say V( r, t) , but thats gets messy

OpenStudy (kainui):

No, that would be wrong like saying V(S,r,t)

OpenStudy (perl):

ok then its just V(r(t))

OpenStudy (kainui):

Maybe you could get fancy and write something like V(S(r(t))) that could be perfectly fine although super weird lol

OpenStudy (perl):

V(r(t)) = g(t)

OpenStudy (perl):

right, lets not complicate it further

OpenStudy (perl):

i just needed to clarify it in my head :)

OpenStudy (perl):

did you finally get the same answer dV/dt = -C * V^(2/3) ?

OpenStudy (kainui):

You can derive some interesting things in tensor calculus by saying things like y(x(y))=y where the outside y is a function and x(y) is the inverse but y by itself is an independent variable. Take the derivative to get: \[\LARGE y'(x(y))*x'(y)=1\] Then you can turn x(y) into just x and get: \[\LARGE y'(x)=\frac{1}{x'(y)}\] which is the same as saying inverse functions have reciprocal slopes, which makes sense since they are just reflected across the line y=x

OpenStudy (perl):

right

OpenStudy (kainui):

Oh I just showed that my answer satisfies the differential equation above so I'm done checking lol. Just sorta playing around idk